OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <7, -4>, v = <-28, 16>

OpenStudy (unklerhaukus):

take the dot product, if it equal zero the vectors are orthogonal

OpenStudy (anonymous):

How do you take the dot product?

OpenStudy (unklerhaukus):

\[\mathbf a=\langle a_1,a_2\rangle\\\mathbf b=\langle b_1,b_2\rangle\]\[\mathbf a\cdot \mathbf b=\langle a_1,a_2\rangle\cdot\langle b_1,b_2\rangle=a_1b_1+a_2b_2\]

OpenStudy (unklerhaukus):

\[\mathbf c=\langle c_1,c_2\rangle\\\mathbf d=\langle d_1,d_2\rangle\] if the vectors are parallel one will be a scaler multiple of the other \[\mathbf c=\langle c_1,c_2\rangle=\langle \alpha d_1,\alpha d_2\rangle=\alpha \langle d_1,d_2\rangle=\alpha\mathbf d\]

OpenStudy (anonymous):

Thank you so much for your help !

OpenStudy (unklerhaukus):

note: the scalar can be negative